OpenStudy (yuki):

There is a particle that moves on a line with the velocity v = 12sqrt(s) where s is the distance from the origin. If s = 1 when t = 0, what is the value of s when t = 1 ?

OpenStudy (anonymous):

i'm pretty sure you typed the question wrong, velocity should be a function of time so it should be v = 12 sqrt (t) integrate the function, we'll call the new function s you get s(t) = 8t^(3/2) + C plug in t = 0, which would make for the function s(0) = 8t^(3/2) + C 1 = C so s(t) = 8t^(3/2) + 1 s(1) = 8*1(3/2) + 1 s(1) = 9

OpenStudy (yuki):

I thought the same thing at first, but I copied it exactly the way it is. The answer has to be 49

OpenStudy (anonymous):

hmm....is there anything else to the question? i'm sorry for asking this, but it's hard to see velocity as a function of distance.. you usually see it as a function of time and i don't know how you're supposed to integrate that then, what would be the new function?

OpenStudy (yuki):

nope, that's all its written. I am thinking that parametric equations is the key. But the question is too specific for me to look it up.

OpenStudy (anonymous):

I'm sorry, i can't help you even if we integrate with respect to s it wouldnt matter because s is 0 so constant would be very small. plugging 0 OR 1 would give you a small number 49 is a huge number... sorry, but good luck

OpenStudy (yuki):

It's all right. Thanks for trying.